# Data-Driven Robust Predictive Control with Interval Matrix Uncertainty Propagation

### 16 Mar 2026 v1

@@ VI Conclusions
A data-driven robust predictive control scheme has been presented, which is able to deal with the presence of unknown-but-bounded process disturbances affecting the unknown system dynamics. The proposed approach provides a viable alternative to the classical pipeline coupling set-membership system identification and model-predictive control. The combined use of interval matrix uncertainty representation and variable-horizon optimal control is instrumental to guarantee both recursive feasibility of the data-driven control scheme and an efficient uncertainty propagation along the prediction horizon. This results into a computationally lightweight control scheme, compared to tube-based MPC approaches. The investigation of alternative data-driven uncertainty representations, and of the associated trade-off between conservatism and computational load of the resulting control scheme, will be the subject of future research.
@@ A-D Proof of Theorem 2
When $N_{k}^{*}=1$, the cost associated with the candidate solution (63) is $\hat{J}_{k+1}=\gamma$, which implies that such a solution is also optimal (as the cost value cannot be less than $\gamma$). Consequently, for all $k>\bar{k}$ one has $J^{*}_{k}=\gamma$ and the optimal solution of problem (34) is the candidate solution (63). Then, due to Assumption 2, the trajectories of the closed-loop system will remain indefinitely inside the terminal set $\mathcal{X}_{f}$. Moreover, since for $k>\bar{k}$ one has $x(k+1)=\mathcal{A}_{K}x(k)+w(k)$, all trajectories will converge to the set $\sum_{i=0}^{+\infty}\mathcal{A}_{K}^{i}\mathcal{W}\subseteq\sum_{i=0}^{+\infty}\mathbb{T}^{i}_{\mathcal{I}_{A_{K}}}(\mathcal{W})=\mathcal{X}_{\infty}$. ∎

### 18 Mar 2026 v2

@@ VI Conclusions
A data-driven robust predictive control scheme has been presented, which is able to deal with the presence of unknown-but-bounded process disturbances affecting the unknown system dynamics. The proposed approach provides a viable alternative to the classical pipeline coupling set-membership system identification and model-predictive control. The combined use of interval matrix uncertainty representation and variable-horizon optimal control is instrumental to guarantee both recursive feasibility of the data-driven control scheme and an efficient uncertainty propagation along the prediction horizon. This translates into a computationally lightweight control scheme, compared to traditional tube-based MPC approaches. The investigation of alternative data-driven uncertainty representations, and of the associated trade-off between conservatism and computational load of the resulting control scheme, will be the subject of future research.
@@ A-D Proof of Theorem 2
When $N_{k}^{*}=1$, the cost associated with the candidate solution (66) is $\hat{J}_{k+1}=\gamma$, which implies that such a solution is also optimal (as the cost value cannot be less than $\gamma$). Consequently, for all $k>\bar{k}$ one has $J^{*}_{k}=\gamma$ and the optimal solution of problem (37) is the candidate solution (66). Then, due to Assumption 2, the trajectories of the closed-loop system will remain indefinitely inside the terminal set $\mathcal{X}_{f}$. Moreover, since for $k>\bar{k}$ one has $x(k+1)=A_{K}x(k)+w(k)$, all trajectories will converge to the set $\sum_{i=0}^{+\infty}\mathcal{A}_{K}^{i}\mathcal{W}\subseteq\sum_{i=0}^{+\infty}\mathbb{T}^{i}_{\mathcal{I}_{A_{K}}}(\mathcal{W})=\mathcal{X}_{\infty}$. ∎